CH/Agen/94/03

Agenda

Charissa Group Meeting

To be held on: Wednesday, 27 April 1994, 11.30 am

To be held at:  Birmingham

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

1.  Minutes of Meeting of 23 February 1994
2.  Matters Arising / Actions
3. Update on NPC matters (DLW)
4.  Progress with MEGHA development

1. Prototype Module

2. Chamber Design

3. Detectors

4. Gas Supply

5.  Progress with electronics and data acquisition
6.  De-briefing on recent ANU running period
1. Chain state excitation functions  (CDJ/DLW)
2. Hybrid tests  (WNC/GJG)
3. ^{28}Si resonance experiment  (GJG/WNC)
4. MEGHA in-beam tests  (DLW/CDJ)
5. Detector/window stocks and shipping  (RPW/DLW)
6. Other operational matters  (DLW/WNC)
7. Plans for July running^{†} (BRF)
8. Proposed spend of remaining grants money^{†}  (BRF)
9. Author list for ANU publications^{†}
10. Publications^{†}
11. Any other business
12. Date of next meeting

W.N. Catford (Secretary) Anzac Day, 25 April 1994

†Paper to be tabled by BRF